Week of 181029

WLCG Operations Call details

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Portal
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to wlcg-scod@cernSPAMNOTNOSPAMPLEASE.ch to allow the SCOD to make sure that the relevant parties have the time to collect the required information, or invite the right people at the meeting.

Best practices for scheduled downtimes

Monday

Attendance:

  • local: Andrea (Storage), Borja (Chair, Monit), Ivan (ATLAS), Julia (WLCG), Maarten (ALICE), Olga (Computing), Remi (Storage), Vincent (Security), Vladimir (LHCb)
  • remote: Christoph (CMS), Di (TRIUMF), Elena (CNAF), Jens (NDGF), John (RAL), Xin (BNL)

Experiments round table:

  • ATLAS reports ( raw view) -
    • Production - 340k / 381k slots
      • HLT farm in GRID production - 95k slots since Sunday.
    • Storage & Transfers - NTR

  • CMS reports ( raw view) -
    • Production activities still restricted to most important campaigns due to shortage of disk space
    • Instabilities of EOS hosted directories still persist
      • Fuse mount very unstable over the weekend: INC:1784940 (recent updates)
      • Websites frequently not reachable GGUS:137768 (ticket is solved, but issue continues to re-appear!)

  • ALICE -
    • NTR

Andrea pointed out LHCb issues seems to be related to IPv6, he suggested to temporarily remove it while investigation is ongoing. Vladimir and Maarten agreed, it was done and temporarily fixed by the end of the meeting.

Sites / Services round table:

  • ASGC:
  • BNL: NTR
  • CNAF:
  • EGI:
  • FNAL: Took downtime last week to enable dual stack on both dCache instances
  • IN2P3:
  • JINR: NTR
  • KISTI:
  • KIT:
    • Downtime for CMS (cmssrm-kit.gridka.de) on last Tuesday for updating dCache, deploying IPv6 and enable high availability for SRM. The downtime had to be extended to 10 am. on Wednesday because some host certificates were not correctly generated and it took us longer to deploy the high availability on the bonded interfaces of the SRM nodes (a novelty among GridKa setups).
      • Only on Friday we fixed an issue for the CMS AAA redirectors, that were not listening on IPv6.
    • We proposed Tuesday, 7th of Nov, to ATLAS to carry out the same tasks in a downtime for them (atlassrm-kit.gridka.de).
  • NDGF: Scheduled downtime for operation on Wednesday, ATLAS will be offline during that time.
  • NL-T1:
  • NRC-KI:
    • Tape system downtime this week for maintenance
  • OSG:
  • PIC:
  • RAL: NTR
  • TRIUMF: New computing cluster (~146k 64b-bit HEPSPEC06) is now in production.

Maarten asked if there are still resources running in the old Computing Centre for ATLAS, Di confirmed some disk resources are still running there but they will be completely moved to the new one.

  • CERN computing services: NTR
  • CERN storage services:
    • FTS: FTS Prod instance for CMS/LHCb upgrade to FTS 3.8.1 tomorrow from 10 to 12:30 CET (OTG:0046529). The upgrade will be transparent to clients.
    • CASTOR: ongoing preparation for the Heavy-Ion data taking. During this period, very limited tape capacity will be given to tape recalls as all the capacity will be dedicated to storing the incoming data.
  • CERN databases:
  • GGUS: NTR
  • Monitoring:
    • Final reports for the September availability sent around
  • MW Officer:
    • An EGI broadcast has been sent to all EGI sites concerning TLS v1.2 support on EGI / WLCG
  • Networks:
  • Security: NTR

AOB:

Edit | Attach | Watch | Print version | History: r20 < r19 < r18 < r17 < r16 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r20 - 2019-05-13 - DavidMason
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback